# Prompts to Proxies: Emulating Human Preferences via a Compact LLM Ensemble

### v1

@@ 5 Conclusion
This work establishes a theoretical foundation for viewing preference alignment as a problem of preference reconstruction. Inspired by revealed preference theory, our approach formalizes alignment as a two-stage process: first, constructing a functional basis of proxy agents using attribute-guided prompting; and then, recovering aggregation weights through supervised learning. In doing so, our work offers the first tractable implementation of pluralistic alignment, along with quantitative metrics to assess both diversity and fidelity.
Beyond its theoretical contributions, this framework offers a practical platform for interdisciplinary research. In the social sciences, it can support survey design, question testing, and nonresponse mitigation, serving as a cost-effective complement to traditional data collection. For researchers in LLM alignment, the modularity of our system provides a controlled setting to evaluate prompt engineering strategies and their influence on agent diversity and downstream performance. More broadly, it lays the groundwork for agentic systems that enable value-sensitive modeling across disciplines. Notably, P2P repositions social survey data not as training material, but as a guiding signal for culturally grounded model alignment.
Appendices
Appendix A Simulation Studies
Simulation Design
Construction of the Ground Truth Data
To construct the ground truth data, we sample a fraction of the endowments as the ground truth and randomly assign weights to these endowments, which represent the proportions of the associated personae in the ground truth population. The ground truth data—percentages of individuals choosing difference options for a given question—are computed using weighted averages of responses provided by the ground-truth agents. The unselected endowments are used to construct the pool of proxy agents.
Regression Method
A.1 Recovering Ground Truth Agents with Constrained Lasso
A.2 Emulating Ground Truth Responses with Proxy Agents
$H_{i}(\mathcal{A})=\frac{-\sum_{k=1}^{K_{i}}p_{ik}\log_{2}p_{ik}}{\log_{2}K_{i}}$ (5)
${H}(\mathcal{A})=\frac{1}{N}\sum_{i=1}^{N}H_{i}(\mathcal{A}).$ (6)
$ECR=\frac{{H}(\mathcal{A}_{\text{proxy}})}{{H}(\mathcal{A}_{\text{gt}})}.$ (7)
Figure 9: Entropy Coverage Ratio Diagnostics for Simulation Study 2.
A.3 Response Variability Matters
Figure 10: Endowment Count per Mode (Sorted by Entropy).
@@ A.3 Response Variability Matters
  3. 3.
Figure 12 offers three snapshots of the simulation runs corresponding to ECRs equaling $0.50$, $1.01$ and $1.67$. When proxy entropy is low compared with ground truth, the proxy agents fail to provide a sufficiently expressive basis to span the aggregate ground-truth preference. As a result, constrained Lasso can only approximate a lower-dimensional projection of the true preference signal. This results in poor prediction performance, as illustrated in the leftmost snapshot. Conversely, when proxy entropy exceeds ground-truth entropy, prediction error begins to increase—albeit at an infinitesimal scale. The subtle deterioration of performance when ECR surpasses 1 could be due to various reasons. A plausible cause is from the data perspective: when the proxy pool becomes too diverse, it may introduce spurious variability that does not align well with the true underlying structure. Put differently, an observed response might amalgamate signals from different latent preference factors, some of which are irrelevant to the aggregate preference in focus. While greater diversity expands the representational capacity of the proxy pool, it may also dilute the signal and induce overfitting, thereby reducing generalization accuracy.
Appendix B Taking P2P to the Wild: Supplementary Experiment Results
B.1 Wave 42
Tracking entropy during active endowment generation
Figure 13 shows the entropy trajectories by question during active endowment generation. Of the tracked questions, 53 experience a clear rising trend in entropy, while 3 see a noticeable fall in entropy at the end of the generation, the other questions have stable entropy values across the update steps. Figure 14 displays the question entropies at the end of active endowment generation.
Figure 13: Entropy trajectories by question during active endowment generation. Trends are classified based on slope and volatility of the entropy trajectory: “rising” if the slope exceeds 0.01 and standard deviation is above 0.02; “falling” if the slope is below –0.01 with sufficient volatility; otherwise labeled “stable”.
Lasso Selection
Figure 15 plots the cross-validated mean squared error (MSE) for each value of $\alpha$ examined during hyperparameter tuning. Among the 30 candidate values, cross-validation identifies the optimal $\alpha$ as $1.17\times 10^{-3}$. Figure 16 shows the number of agents selected by constrained lasso by mode entropy, together with the number of agents selected from each template/strategy. Mixed modes contribute the largest share to the ensemble with 20 proxy agents, followed by question-patching modes with 14. Thematic modes come next, contributing 5 agents, alongside 5 from the survey mode. The remaining 1 proxy is selected from the core mode.
Figure 15: Cross-validated MSE by alpha ($\alpha$), the penalization parameter, for constrained lasso. The best alpha selected by CV is $1.17\times 10^{-3}$.
Figure 16: Number of agents selected by constrained lasso by mode entropy. Mixed modes contribute 20 proxies to the ensemble, followed by question patching modes contributing 14. The thematic modes come third, contributing 5 proxies. Five proxy agents are selected from the survey mode, while the rest 1 is from the core mode.
Elastic Net Results
In the main paper, we show prediction results using constrained lasso. Here, we provide the results for constrained elastic net (Figures 17 and 18). Overall, the prediction results are similar those using constrained lasso.
Figure 17: Cross-validated MSE by alpha, the overall penalization parameter, and L1 ratio, the penalization weight for the L1 term, for constrained elastic net. The lowest cross-validated MSE is achieved at $\alpha=2.59\times 10^{-3}$, with l1 ratio set at $0.60$.
Figure 18: Prediction performance of the agent ensemble selected by constrained elastic net. Results are similar to those using constrained lasso.
B.2 Panel Study Across ATP Waves
To obtain a panoramic view of P2P’s performance empirically, we conduct a preliminary panel study applying P2P to a collection of 14 waves from the American Trends Panel (ATP). Each wave is repeated three times to assess the consistency and robustness of our method under varying random seeds and endowment draws. This subsection provides an extended empirical perspective on the real-world applicability of our approach and sheds light on variability patterns that may not arise in controlled simulation environments. The ATP waves used in this panel study are: W26, W27, W29, W32, W34, W36, W41, W42, W45, W49, W50, W54, W82, and W92.
Figure 19 illustrates the test MSEs for the 42 experiments organized by the number of survey questions per wave. As shown, within-wave performance is generally stable, with most waves exhibiting low variance in test MSE across repeats. However, test MSE varies more significantly across waves. The performance distribution is summarized below:
  • 8/14 waves achieve test MSE consistently below 0.015.
  • 4/14 waves fluctuate around 0.015.
  • 2/14 waves have higher error ranges (0.020–0.025).
These preliminary results reveal no clear relationship between test MSE and the number of survey questions, suggesting that wave-specific factors, such as survey content, attribute coverage, or question formulation, may play a larger role in shaping performance.
Figure 19: Test performance across ATP waves. Each point represents a wave-specific run, colored by wave identity. The horizontal dashed lines mark two threshold levels: 0.015 and 0.020, used to classify performance regimes.
To illustrate the interaction between question-level entropy dynamics and predictive performance, we include three snapshots of the panel study in Figure 20, showcasing the differences in question entropy trajectories during active endowment generation and prediction performance.
Figure 20: Representative snapshots for three ATP waves with different test-time MSE. Top: the worst-performing wave. Middle: a moderately performing wave. Bottom: the best-performing wave. Left: entropy trajectories classified as rising (green), falling (red), or stable (gray). Center/right: Lasso regression fits for aggregate response prediction. While rising entropy often coincides with improved fit, some waves achieve strong performance with limited entropy change, suggesting additional contributing factors beyond diversity expansion.
Interestingly, the wave with the lowest test MSE (bottom row) exhibits relatively modest entropy change, while the poorest-performing wave (top row) shows mostly flat entropy evolutions. This suggests that while entropy dynamics can be indicative of diversity expansion, strong initial endowment coverage or other wave-specific factors (e.g., question content or noise) may also influence alignment outcomes. Entropy trends are a useful diagnostic but should be interpreted in the context of wave-specific factors.
The patterns observed in the panel study warrant deeper analysis, which is likely to inform refinements that make P2P more stable and effective across settings. One especially promising direction is to extend the entropy-based active endowment generation strategy to support multi-objective criteria, allowing the system to balance diversity, coverage, and alignment more systematically.
Another promising extension is to include a module that directly measures question semantics. Based on our preference reconstruction theory, survey questions define the response space $\mathcal{R}$. Question entropy serves as a key signal for underlying diversity when conditioning on a fixed question. However, an equally important source of variation lies in the diversity across questions themselves. This includes differences in topic coverage, difficulty, and framing—factors that shape the expressiveness and completeness of the response space. Ideally, the question set should span a semantically broad topic space so that variations in response entropy meaningfully reflect differences in latent preferences. A dedicated semantic module would enable content-aware data-splitting strategies and support more rigorous generalization analysis across question types. This is an important direction for future development of P2P.
Due to budget constraints, our empirical evaluation relies on American Trends Panel (ATP) waves, which are topical in nature and tend to focus on a narrow set of issues per wave. While this allows for controlled testing, it limits the diversity of question types and themes. A more expansive and promising application domain for P2P is large-scale general social surveys, such as the General Social Survey (GSS) or World Values Survey, which span a wide range of topics and include high-quality, theoretically grounded questions relevant across disciplines. Testing P2P in such settings would provide a more comprehensive benchmark for its alignment and coverage capabilities, and yield better insights into its generalizability across domains of social inquiry. Moreover, the agent ensembles generated from general social surveys could, in turn, be used to provide feedback on smaller topical surveys to improve survey design and assess content quality.
Appendix C Methodological and Experimental Details
C.1 Prompts
C.1.1 Attribute Learner (AttributeLearner)
@@
You are an intelligent research assistant trained to analyze individual survey questions and infer which human attributes might influence how different people respond.
You are given a single survey question. Your task is to propose a list of relevant human attributes—such as demographics, beliefs, values, personality traits, or ideological leanings—that are likely to shape responses to this question.
Focus on underlying factors that would cause meaningful variation in answers across different types of people. Avoid generic or overly broad attributes.
Respond **only** with a Python-style list of double-quoted strings. Do not include any explanation, headers, or prose before or after the list.
**Example output format:** ["religious affiliation", "political ideology", "trust in government"]
@@
You are an intelligent research assistant trained to analyze survey questions and infer which human attributes might influence how individuals respond.
You are given a set of training-only survey questions. Your task is to propose a list of relevant human attributes—such as demographics, beliefs, values, personality traits, or ideological leanings—that are likely to affect responses to these questions.
Carefully analyze the content and framing of the questions. Identify underlying factors that might shape how different people respond. Focus on attributes that are salient, discriminative, and potentially variable across respondents.
Respond **only** with a Python-style list of double-quoted strings. Do not include any explanation, headers, or prose before or after the list.
**Example output format:** ["attribute1", "attribute2", "attribute3", "attribute4"]
C.1.2 Endowment Model (EndowmentModel)
To generate endowments from attributes:
You are an expert assistant trained to generate realistic, diverse, and demographically plausible personas for social science surveys.
Each persona should include:
‘eid‘: a short, lowercase, variable-safe identifier that encodes key traits (e.g., urban_liberal_30s_female). No punctuation or spaces.
‘endow_text‘: a natural language description of the persona (1-2 sentences), written as if describing a survey respondent.
Instructions:
Represent a wide range of age, gender, race, education, region, and political ideology.
Avoid repetition of phrasing or demographic combinations across personas.
Do not include explanations or formatting outside of the JSON array.
C.1.3 Survey Conductor (SurveyConductor)
@@
You are completing a survey.
Your answer should reflect the person described in the profile above, using their preferneces, beliefs and experiences.
Respond with only the final answer string, not the code or label in brackets.
Do not include any reasoning, explanation, or commentary.
Do not preface your answer with phrases like ’I would choose’.
Just return the answer text exactly as it appears in the options.
C.2 Active Endowment Generation

### v2

@@ 5 Conclusion
This work formalizes preference alignment as a problem of preference reconstruction and operationalizes it as a two-stage process: first, constructing a functional basis of proxy agents; and then, recovering aggregation weights through supervised learning. We introduce P2P, a modular alignment system that implements preference reconstruction, combining entropy-based adaptive sampling with regression-based aggregation. Our simulation results corroborate P2P’s design choices. Empirical performance shows P2P can recover preference patterns observed in real human survey data while remaining competitive against SFT-aligned models despite using significantly less data. Future work can extend P2P to accommodate freeform textual output and aggregation, analyze its use as a benchmark for model steerability, devise re-anchoring strategies for non-stationary preference, and incorporate semantic and psychometric techniques for improved robustness and generalization (Appendix F). Our work offers a first step towards these endeavors and a practical platform for interdisciplinary research.
Acknowledgments
This work was partially supported by Google Cloud free trial credits and by discretionary research funds and compute resources from the GLOW.AI group, led by Prof. Bryan Low. The views and interpretations expressed in this paper are solely those of the authors.
Impact Statement
This work uses publicly available, anonymized survey datasets with sensitive identifiable information removed. While the large language model (LLM) backends have been trained to align with human values, they may still propagate social biases and stereotypes, and some generated endowment profiles may contain opinionated content. Our goal is methodological—studying preference reconstruction at the aggregate level rather than individual personalization—which reduces profiling risks. We see potential for positive impact on pluralistic alignment and principled use of AI in social science research, but stress that the agent ensemble is an auxiliary tool for understanding human choice patterns and should not replace humans in critical policy-making scenarios.
Appendix A Modeling Preferences via Functional Bases: Proofs
Proposition A.0 (Existence of a representative agent).
Let $\widetilde{\mathcal{P}}$ be a latent preference space (regardless of human or model), with an individual preference denoted by $\tilde{p}\in\widetilde{\mathcal{P}}$. Let $g$ be the map from $\widetilde{\mathcal{P}}$ to the observed response space $\widetilde{\mathcal{R}}\subset\mathbb{R}^{n}$. Then the following two statements are equivalent:
@@ Appendix A Modeling Preferences via Functional Bases: Proofs
    For any aggregate response $\tilde{r}_{\mathrm{pop}}=\sum_{i=1}^{I}w_{i}\tilde{r}_{i}$ with $\{w_{i}\}_{i=1}^{I}$ convex weights and $\tilde{r}_{i}=g(\tilde{p}_{i})$, there exists $\tilde{p}_{\mathrm{pop}}\in\widetilde{\mathcal{P}}$ such that $g(\tilde{p}_{\mathrm{pop}})=\tilde{r}_{\mathrm{pop}}$.
Proof.
First, we prove $(1)\implies(2)$.
[…]
By the definition of image, it follows that there exists $\tilde{p}_{\mathrm{pop}}\in\widetilde{\mathcal{P}}$ such that
$g(\tilde{p}_{\mathrm{pop}})=\sum_{i}^{I}w_{i}\tilde{r}_{i}\triangleq\tilde{r}_{\mathrm{pop}}.$
Claim: $(2)\implies(1)$.
[…]
Remark A.1 (Sufficient condition for representative agent).
If the decision-making mechanism $g$ maps a preference into a convex subset of a probability simplex $\Delta^{n-1}$, then any aggregate of individual responses is realizable by a representative agent.
Proof.
A convex subset of a probability simplex is by definition convex. The result follows from the theorem. ∎
The implications of this lemma are two-fold:
  1. 1.
[…]
  2. 2.
[…]
See 2.1
Definition A.2.
[…]
See 2.2
Definition A.3.
[…]
Proposition A.4.
Human population preference is learnable in the revealed sense if and only if individual human preference is learnable in the revealed sense.
Proof.
The “only if” part follows directly from $h(\mathcal{P})\subset\mathrm{conv}(h(\mathcal{P}))$. Here we focus on the “if” part.
By definition, a human population preference is defined as
$r_{\mathrm{pop}}=\sum_{i=1}^{I}w_{i}h(p_{i}).$
Since individual human preference is learnable in a revealed sense, i.e., $)h(\mathcal{P})\subset\mathrm{conv}(f(\mathcal{\widehat{P}}$)), for any $p_{i}$, there exists $\{w_{ik},\hat{p}_{ik}\}_{k=1}^{K_{i}}$ such that
$h(p_{i})=\sum_{k=1}^{K_{i}}w_{ik}f(\hat{p}_{ik}).$
Plug it back into the definition of human population preference and we get
[…]
The last step follows from re-indexing. Therefore, $)\mathrm{conv}(h(\mathcal{P}))\subset\mathrm{conv}(f(\mathcal{\widehat{P}}$)). ∎
See 2.3
Proof.
By Proposition A.0, if human population preference can be learned by a representative LLM agent, then
[…]
For any $r_{\mathrm{pop}}\in\mathrm{conv}\left(h(\mathcal{P})\right)$, it follows from the definition of human population preference that
$r_{\mathrm{pop}}=\sum_{i=1}^{I}w_{i}h(p_{i}):=\sum_{i=1}^{I}w_{i}r_{i}.$
[…]
It follows that
$r_{\mathrm{pop}}=\sum_{i=1}^{I}w_{i}f(\hat{p}_{i})\in\mathrm{span}\left\{f(\hat{p}_{i})\right\}_{i=1}^{I}.$
In words, we can construct an ensemble where each agent represents in preference a ground-truth individual in the human population (demographic conditioning).
NB: Here we implicitly assume that the human population consists of more than one individual with distinct preferences. If the population consists of only 1 individual, then the representative agent and the ensemble coincide. ∎
See 2.4
Proof.
[…]
$).\mathrm{conv}(h(\mathcal{P}))\subset\mathrm{conv}(f(\widehat{\mathcal{P}}$)). (9)
[…]
$r_{\mathrm{pop}}=\sum_{j=1}^{J}\tilde{w}_{j}f(\hat{p}_{j}).$
In other words, human population preference can be learned using an ensemble. ∎
See 2.5
Proof.
[…]
$r_{\mathrm{pop}}=\sum_{i=1}^{I}w_{i}r_{i}=\sum_{i=1}^{I}w_{i}f(\hat{p}_{i}).$ (10)
However, since $r_{1}\in h(\mathcal{P})\subset\mathrm{conv}(h(\mathcal{P}))$, we know there exists a finite set of $\{r^{*}_{k}\}_{k=1}^{K}$ with ${r}_{k}^{*}\in h(\mathcal{P})$ and convex weights $\{\lambda_{k}\}_{k=1}^{K}$ such that
$r_{1}=\sum_{k=1}^{K}\lambda_{k}{r}_{k}^{*}.$
Plug this into (10) and we get
$r_{\mathrm{pop}}=\sum_{k=1}^{K}w_{1}\lambda_{k}{r}_{k}^{*}+\sum_{i=2}^{I}w_{i}r_{i}.$
[…]
In conclusion,
$r_{\mathrm{pop}}\in\mathrm{span}\{f(\hat{p}_{1}^{*}),\dots,f(\hat{p}_{K}^{*}),f(\hat{p}_{2}),\dots,f(\hat{p}_{I})\}.$
Appendix B Prompts to Proxies (P2P): Implementation Details
B.1 Attribute Bank
For the preset attribute bank (Figure 4), we define three primary templates. The core template contains general demographic descriptors and serves as the default baseline. The thematic template includes topic-specific modes such as economics, politics, culture and history. The theoretical template groups attributes derived from canonical frameworks such as Maslow’s hierarchy of needs (Maslow, 1943) and the Big Five personality traits (John and Srivastava, 1999). These templates and modes are not fixed: they serve as a structured starting point but are fully extensible. Users can define new templates, introduce domain-specific modes, or modify existing attribute sets to suit different experimental contexts. By sampling and varying attributes from this bank, we create prompts that instantiate proxy agents at different locations of the latent model preference space.
Figure 4: Attribute Bank and Endowment Generation Workflow. The preset attribute bank is organized hierarchically by template (core, thematic, theoretical), mode (e.g., economics, politics, maslow, big five), and attribute (e.g., race, gender, ethnicity, urban/rural residency, risk aversion). Endowment is generated by a specialized LLM (EndowmentModel), which takes drawn attributes as input and outputs a profile with instantiation of these attributes.
B.2 Model Prompts
B.2.1 Attribute Learner (AttributeLearner)
@@
Attribute Learner: Learning Attributes from A Single Question You are an intelligent research assistant trained to analyze individual survey questions and infer which human attributes might influence how different people respond. You are given a single survey question. Your task is to propose a list of relevant human attributes—such as demographics, beliefs, values, personality traits, or ideological leanings—that are likely to shape responses to this question. Focus on underlying factors that would cause meaningful variation in answers across different types of people. Avoid generic or overly broad attributes. Respond **only** with a Python-style list of double-quoted strings. Do not include any explanation, headers, or prose before or after the list. **Example output format:** ["religious affiliation", "political ideology", "trust in government"]
@@
Attribute Learner: Learning Attributes from A Set of Questions You are an intelligent research assistant trained to analyze survey questions and infer which human attributes might influence how individuals respond. You are given a set of training-only survey questions. Your task is to propose a list of relevant human attributes—such as demographics, beliefs, values, personality traits, or ideological leanings—that are likely to affect responses to these questions. Carefully analyze the content and framing of the questions. Identify underlying factors that might shape how different people respond. Focus on attributes that are salient, discriminative, and potentially variable across respondents. Respond **only** with a Python-style list of double-quoted strings. Do not include any explanation, headers, or prose before or after the list. **Example output format:** ["attribute1", "attribute2", "attribute3", "attribute4"]
B.2.2 Endowment Model (EndowmentModel)
System prompt:
Endowment Model: System Prompt You are an expert assistant trained to generate realistic, diverse, and demographically plausible personas for social science surveys. Each persona should include: - ‘eid‘: a short, lowercase, variable-safe identifier that encodes key traits (e.g., urban_liberal_30s_female). No punctuation or spaces. - ‘endow_text‘: a natural language description of the persona (1-2 sentences), written as if describing a survey respondent. Instructions: - Represent a wide range of age, gender, race, education, region, and political ideology. - Avoid repetition of phrasing or demographic combinations across personas. - Do not include explanations or formatting outside of the JSON array.
To generate endowments based on attributes:
Endowment Model: Endowment Generation Based on Attributes Generate {n} diverse persona(s) that vary meaningfully along the following attributes: {attr_string}. Each persona should reflect a distinct combination or value expression of these traits. Return a JSON array of dictionaries, each with: - ‘eid‘: a short, lowercase, variable-safe identifier - ‘endow_text‘: a brief natural-language description of the persona
To generate endowments based on survey topics (also used for Vanilla baseline):
Endowment Model: Endowment Generation Based on A Survey Topic Generate {n} diverse personas for a survey experiment. Each persona must include: - eid: short, lowercase identifier - endow_text: a short natural-language description The survey topic is: {topic} Ensure diversity across demographics. Return the result as a JSON array.
B.2.3 Survey Conductor (SurveyConductor)
@@
Survey Conductor: Answering A Survey Question You are completing a survey. Your answer should reflect the person described in the profile above, using their preferneces, beliefs and experiences. Respond with only the final answer string, not the code or label in brackets. Do not include any reasoning, explanation, or commentary. Do not preface your answer with phrases like ’I would choose’. Just return the answer text exactly as it appears in the options.
B.3 Active Endowment Generation